At what altitude must a military plane flying horizontally at 1.75m/s be in order to deliver relief goods to typhoon yolanda survivors? The survivors are in an evacuation center 1750m away from where the plane is.

1
Expert's answer
2021-11-08T11:52:42-0500

The time to pass 1750 m at 1.75 m/s:


"t=d\/v."

The height of free fall that corresponds to this height:


"H=gt^2/2=gd^2/(2v^2)=4.9·10^6\text{ m}."